OpenStudy (anonymous):

1. A sample consisting of 1.0 mol CaCO3 (s) was heated to 800oC when it is decomposed. The heating was carried out in a container fitted with a piston that was initially resting on the solid. Calculate the work done during complete decomposition at 1.0 atm. What work would be done if instead of having a piston, the container was open to the atmosphere?

OpenStudy (anonymous):

For the first part: Vf >> Vi so: Vf = nRT/P(external) This gives us: w = -P(ex) *(nRT/P(ex) w = - nRT w= -(1.0 mol)(8.314*10^-3 kj/K mole)(1073.15 K) w=-8.92 kj

OpenStudy (anonymous):

Also the second part would also be the same (w=-8.92) because the piston is allowed to move in the first part. So, in the second, it is the same thing without the piston. In both cases the container is open to the atmosphere. So, work is going to be the same in both situations.